OpenStudy (anonymous):

An elevator with mass 3000kg hangs still on a floor that's 5m above a spring(k=1,5*10^5). The cable of the elevator snaps and it falls with constant friction (4500N)

OpenStudy (theeric):

Hello! What is the question, here? It seems some has bee left out!

OpenStudy (anonymous):

1)calculate the speed of the elevator before it hits the spring

OpenStudy (anonymous):

2) The distance the spring will get pressed

OpenStudy (anonymous):

3) The height the spring pushes the elevator up

OpenStudy (anonymous):

it's alot..

OpenStudy (theeric):

Okay! Any idea how to start? Forewarning, I'll have to go in a little bit.

OpenStudy (anonymous):

no idea, but i have an exam tomorrow so i'll have to check it tomorrow xD

OpenStudy (theeric):

So, how should we get started? Falling object because of gravity, opposed by friction... What's the net force?

OpenStudy (anonymous):

it's own weight?

OpenStudy (theeric):

That, and the friction! This elevator is sqeezed in the walls, or maybe is on a pulley that has friction. Either way, the force of friction, \(F_f=4,500\ [N]\), as the problem states. And the force of weight is \(F_\text w=m_\text{elevator}\ g\) where \(g\) is the acceleration to gravity.

OpenStudy (theeric):

\(g\) is a negative number, for the record. What do you use, \(-9.8\ [m/s^2]\)?

OpenStudy (anonymous):

yea

OpenStudy (theeric):

And the net force is \(F_\text{net}=F_\text w +F_f\). So, shall we use substitution? \(F_\text{net}=m_\text{elevator}\ g+F_f=(3,000\ [kg])\ (-9.81\ [m/s^2])+4,500\ [N]\)

OpenStudy (theeric):

\(F_\text{net}=(3,000\ [kg])\ (-9.8\ [m/s^2])+4,500\ [N]\), sorry.

OpenStudy (theeric):

I got \(F_\text{net}=-24,900\ [N]\), but you should probably double check me...

OpenStudy (anonymous):

i would, but i'll have to check tomorrow :P , doing this while studying for my exam :D

OpenStudy (theeric):

Now, it has a net force on it, because \(F_\text{net}\neq0\). So that acceleration gives it it's speed that we need to find after 5 meters. Okay, good luck studying! \(F_\text{net}=m_\text{elevator}\ a\), so \(a=\dfrac{F_\text{net}}{m_\text{elevator}}\) And the velocity? \(v_f^2=v_i^2+2\ a\ d\implies v_f=\sqrt{v_i^2+2\ a\ d\ }\).

OpenStudy (theeric):

Then the elevator has a kinetic energy of \(E_k=\frac{1}{2}m\ v^2\). To stop it, all that energy must be converted to (spring) potential energy as the elevator's momentum does work on the spring. The work on the spring is given by \(W=\frac{1}{2}k\ x^2\) I think. I have to go now.. But \(E_k=W\), and so you can solve for \(x\) to see how far the spring was compressed.

OpenStudy (anonymous):

Thanks theEric ;)

OpenStudy (theeric):

Oh wait, I neglected friction.. Maybe I shouldn't have... On the way up, the spring gives it velocity, but gravity accelerates it downwards. If you can, find out if there is still friction when it is on the spring! Good luck!

OpenStudy (theeric):

\(\sf\Large\color{orange}{Compressing\ the\ spring.}\) So the kinetic energy that the elevator has when it hits the spring will be completely converted to spring potential energy and energies associated with friction (heat, sound, transferred kinetic...). So \(E_\text{k, before}=E_\text{P, after}+E_{f\text{, after}}\) \(E_K=\frac{1}{2}m\ v^2\) \(E_P= \frac{1}{2}k\ y^2\) \(E_f=y\ F_f\) So you can substitute them in to the above equation and solve for \(y\). It shouldn't be too tricky.

OpenStudy (theeric):

Oh, there is also the work done by the weight as the spring is compressed. So \(E_\text{k, before}+E_w=E_\text{P, after}+E_{f\text{, after}}\) \(E_w=m\ g\ y\), which makes it at least a little harder to solve for \(y\).

OpenStudy (theeric):

\(\sf\Large\color{lime}{Decompressing\ the\ spring.}\) Now the energy is this: \(E_{P\text{, spring}} +E_w+E_f=E_K\) How high does the elevator go? Well, \(E_K=0\) when the elevator has stopped going up! Notice that, because of the direction of the forces, \(E_w\) and \(E_f\) will be negative. So set \(E_K\) equal to \(0\), and solve for \(y\) again.

OpenStudy (theeric):

I think that should be right... But I'm not 100% sure I got all of it right. But I hope it was valuable input, anyway! Best of luck!